Determine if the given sets are closed or not closed and why.
1. 10/2=5
2. 50/2=25
3. 45/15=3

Answer by ikleyn(52785)

.
Determine if the given sets are closed or not closed and why.
1. 10/2=5
2. 50/2=25
3. 45/15=3
~~~~~~~~~~~~~~~~~~~~


What is the meaning of this post ?

As I see it, it is simply soup of words, inedible, i.e. with no sense.


What are "the given sets" ? - - - There is no any set given or defined in the post.

What does it mean this term "closed or not closed" ? - - - The meaning is not defined.


Again, the meaning of this post is 0 (= zero).